OpenStudy (anonymous):

solve: 7x+3(4x-1)=-79

OpenStudy (tkhunny):

What's your plan? Collect the 'x's. Distributive Property?

OpenStudy (anonymous):

to solve for x

OpenStudy (tkhunny):

Excellent. What's stopping you?

OpenStudy (anonymous):

adding the like terms

OpenStudy (anonymous):

just the process

OpenStudy (tkhunny):

Work on this part, first: 3(4x-1)

OpenStudy (anonymous):

okaii

OpenStudy (anonymous):

12x-3

OpenStudy (anonymous):

7x+3(4x-1)=-79 7x+12x -3=-79 19x=-79+3 19x=-76 x=-79/19 x=-4.1

OpenStudy (tkhunny):

Perfect. Now relace that part of the equation. 7x+3(4x-1)=-79 7x + 12x - 3 = -79 Now, the "like" terms are more easily identified. What's next?

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

nothing thank you sooo much

OpenStudy (tkhunny):

Be careful not to repeat Chica's error. Did you see it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my error, where is that?

OpenStudy (anonymous):

well its not an error its just the way u wrote part of it i kind of got confused but then i got it

OpenStudy (anonymous):

okay Kayla

OpenStudy (anonymous):

wat?

OpenStudy (tkhunny):

@ Chica No, it's an error. You had -76, but then it magically turnd back into -79.

OpenStudy (anonymous):

okay i saw it, thanks for identifying the problem

OpenStudy (anonymous):

is there any way i can edit it?

OpenStudy (anonymous):

14x+5(4-2x)=20

OpenStudy (tkhunny):

I know of no way to edit, other than [Copy], [Delete], [Repost]. Don't worry about it. It's in the discussion and is unlikely to lead to any confusion.

OpenStudy (tkhunny):

Okay, @KaylaStaff, this is EXACTLY like the last one. I recommend very strongly that you do the first demonstration.

OpenStudy (anonymous):

i tried but im stuck @tkhunny

OpenStudy (tkhunny):

You have to do better than that. Take a deep breath. Think about how we did the last one. Take another deep breath. Apply what you learned from the last one to this one. Distributive Property Like Terms etc. Go!

OpenStudy (anonymous):